# An $\ell^{1}$-Plug-and-Play Approach for Magnetic Particle Imaging Using a Zero Shot Denoiser with Validation on the 3D Open MPI Dataset

### v1

@@ 2.3 The Proposed Plug-and-Play Approach with a Zero-Shot Denoiser
Choosing auxiliary variables $u_{1}$, $u_{2}$ and $u_{3}$ we decouple the quantities in Eq. (3), in particular, we choose the following equivalent constrained minimization problem in consensus form:
@@
In each iteration $k$, the term in Eq. (6) is in charge of the reconstruction whereas the term in Eq. (7) performs regularized Gaussian denoising with positivity constraint. Finally, the term in Eq. (8) is the proximal mapping of the $L^{1}$ norm, i.e., it results in a soft-thresholding.
@@
The idea behind the PnP approach is to substitute the Gaussian denoising in Eq. (7) by any other denoiser with positive output and w.r.t. noise level $\lambda/\mu_{k}$, yielding the following algorithm:
$u_{1}^{k+1}$ $=\arg\min_{u_{1}}\lVert f-Au\rVert_{2}^{2}+\mu_{k}\left\lVert u_{1}-\frac{u_{2}^{k}+u_{3}^{k}}{2}\right\rVert_{2}^{2}$ (9)
@@
The choice of a suitable denoiser in Eq. (10) is key because the performance of the algorithm significantly depends on it. In Computer Vision, machine learning based 2D denoisers have gained a lot of interest in recent years. They are typically trained on large datasets, and possess a higher generalization ability than smaller models.
The denoiser we employ in this work is the benchmark deep denoiser prior [70]. It is based on a deep CNN architecture (DRUNet) which combines a U-Net [49] and ResNet [28]. The denoiser is publicly available and trained on a large dataset that combines various datasets including BSD [11], Waterloo Exploration Database [44], DIV2K [1] and Flick2K [43]. We do not retrain the denoiser on MPI or MPI related data, but employ it as a zero-shot denoiser (in the sense that it has not seen the domain before.)
We perform the denoising of the 3D target in a slicewise fashion strategy [4], i.e., at each iteration $k$ we consider the 2D slices perpendicular to an axis and perform the denoising of each 2D slice with the 2D denoiser. This slicewise denoising is perform at $k$ along each axis and the three resulting volumes are averaged.
The overall algorithm is given in Algorithm 1.
Algorithm 1 The proposed $\ell^{1}$-PnP algorithm.
Input: scan data $f$ and system matrix $A$, parameters $\alpha$, $\mathrm{toll}_{\sigma}$, $\mu_{0}$.
@@
$u_{2}^{0},u_{3}^{0}\leftarrow 0$;
$k\leftarrow 0$;
while $\sigma_{\mathrm{res}}>\mathrm{toll}_{\sigma}$ do
     $u_{1}^{k+1}\leftarrow$ConjGrad$\left(A^{\ast}A+\mu_{k}\mathrm{Id}\,,A^{\ast}f+\mu_{k}\frac{u_{2}^{k}+u_{3}^{k}}{2}\right)$;
     $\sigma_{k}\leftarrow\sqrt{\mathrm{Var}(u_{1}^{k+1})}$;
     if $k=0$ then
         $\lambda\leftarrow\mu_{0}\cdot\sigma_{0}^{2}$
     end if
     $u_{2}^{k+1}\leftarrow$Slicewise-Denoise$\left(u_{1}^{k+1}\,,\sigma_{k}\right)$;
     $u_{3}^{k+1}\leftarrow$Soft-Threshold$\left(u_{1}^{k+1}\,,\frac{\alpha}{\mu_{k}}\right)$;
     $\sigma_{\mathrm{res}}\leftarrow|\sigma_{k}-\sigma_{k-1}\rvert$;
     $\mu_{k+1}\leftarrow\lambda/\sigma_{k}^{2}$;
     $k\leftarrow k+1$;
end while
return $u_{2}^{k+1}$
@@ 2.4 Parameter Selection Strategies
The $\ell^{1}$-PnP approach proposed depends on the (hyper) parameters $\mu_{k}$, $\lambda$ and $\alpha$ as well as on the number of iterations $k$ of the scheme (stopping criterium). (Note that the $\sigma_{k}$ are not free parameters, but related via $\sigma_{k}=\sqrt{\lambda/\mu_{k}}$.) Typically, in PnP approaches (e.g., [70]) the corresponding parameters are set fixed, in particular the noise related parameter $\sigma_{k}$ fed into the denoiser is set fixed for each iteration $k$ for a fixed number of iterations. We now propose more adaptive ways to choose the above parameters.
@@
where $\bar{u}_{1}^{k}$ is the mean of the vector $u_{1}^{k}$. The value $\tilde{\sigma}_{k}^{2}$ is an upper bound on the noise level (and potentially can be seen as a very coarse estimate which empirically turns out to be sufficient for our purposes; another interpretation can be as a normalization.) Then, we set the $\mu_{k}$ by
@@
To determine $\lambda=\tilde{\lambda},$ we use the relation $\tilde{\lambda}=\tilde{\mu}_{0}\tilde{\sigma}_{0}^{2},$ where we define $\tilde{\sigma}_{0}^{2}$ via Eq. (12), and $\tilde{\mu}_{0}$ by the following observation: initializing $u_{2}^{0}=u_{3}^{0}=0$ in Alg. 1, $u_{1}^{0}$ is the solution of the Tikhonov regularization problem (cf. Eq. (9) and Eq. (2) with $\mathcal{R}(u)=\lVert u\rVert_{2}^{2}$). Consequently, one can set the parameter $\mu_{0}$ to
@@
where $\tilde{\lambda}_{\mathrm{Tik}}$ denotes an estimate on the parameter $\lambda_{\mathrm{Tik}}$ of the corresponding Tikhonov problem. Viewing Tikhonov regularization as acting on the singular values of the system matrix $A$ we relate $\tilde{\lambda}_{\mathrm{Tik}}$ to the singular values of $A;$ concretely, we here consider (potentially rough estimates) of the smallest singular value $\sigma_{\mathrm{min}}$ and choose $\tilde{\mu}_{0}$ according to the simple criterion $\tilde{\mu}_{0}=\tilde{\lambda}_{\mathrm{Tik}}=100\,\sigma_{\mathrm{min}}^{2}$ to reduce the impact of the small singular values.
@@
In each iteration a soft-thresholding (proximal mapping of the $\ell^{1}$ norm) of parameter $\alpha/\mu_{k}$ is performed (cf. Eq. (11)), and thus all values of magnitude below $\alpha/\mu_{k}$ are set to zero. So, if $c_{\mathrm{rec}}$ is the maximum concentration of the reconstruction, it is necessary that $\alpha<c_{\mathrm{rec}}\cdot\mu_{0}$ not to get the all zeros reconstruction. Therefore, it is reasonable to choose $\alpha$ as some percentage of $c_{\mathrm{rec}}\mu_{0}$. In the Open MPI data set $c_{\mathrm{rec}}\leq 1$, hence $\alpha$ can be though as a percentage of $\mu_{0}$. We refer the interested reader to the Section S4 of the supplementary material for the details. In this work we have set $\alpha=0.005\mu_{0}$ for all experiments.
Stopping Criterion
We propose to consider the increments of the estimated $\tilde{\sigma}_{k}$ in (12), and to stop the iteration when they fall under a fixed tolerance level $\mathrm{tol}_{\sigma},$ i.e., when $\sigma_{\mathrm{res}}^{k}=\lvert\ \tilde{\sigma}_{k}-\tilde{\sigma}_{k-1}\rvert<\mathrm{tol}_{\sigma}$. In this work we have set $\mathrm{tol}_{\sigma}=10^{-4}$ for all experiments.
@@ 3 Results
In this section we quantitatively and qualitatively evaluate the $\ell^{1}$-PnP algorithm proposed in Section 2.3 on the OpenMPI Dataset [39]. Qualitative comparision using the well established quality measures PSNR and SSIM is possible by inferring the groundtruth from a CAD model; details are given in Section 3.1. We compare the proposed $\ell^{1}$-PnP scheme with the Deep Image Prior (DIP) and with the standard Tikhonov $\ell^{2}$-regularization as a baseline; these methods are discribed in Section 3.2. The evaluation is conducted in Section 3.3. In particular, in Experiment 1 we compare the results obtained with the DIP and Tikhonov regularization on the 3D Open MPI Dataset using typical preprocessing. In Experiment 2 we investigate the influence of increasingly lower levels of preprocessing. Finally, in Experiment 3 we show the results obtained with $\ell^{1}$-PnP approach without preprocessing using our parameter selection and stopping strategy (and without relying on a SVD).
The preprocessing was implemented in Matlab and the $\ell^{1}$-PnP algorithm in Python 3.9, using Numpy and PyTorch. The experiments were performed on a workstation with 13th Gen Intel(R) Core (TM) i9-13900KS, 128 GB of RAM, an NVIDIA RTX A6000 GPU and Windows 11 Pro.
@@ 3.1 Image Quality Measures
In order to quantitatively assess the reconstruction quality, the Open MPI Dataset provides CAD models of the phantoms. This allows to compare the reconstructions with the corresponding ground truth implied by the CAD model. We use both Peak Signal-to-Noise Ratio (PSNR) and the Structural Similarity Index Measure (SSIM) as measures. For details on the employed implementation, we refer to Sec. S2 in the supplementary material.
@@
In what follows we always display $\mathrm{PSNR}_{\mathrm{max}}$ and $\mathrm{SSIM}_{\mathrm{max}}$ but stick to the notation PSNR and SSIM for simplicity.
@@ 3.2 Methods for Comparison
The inversion of the system matrix is performed considering the minimization problem in Eq. (2) with $\mathcal{R}(u)=\lVert u\rVert_{2}^{2}$; the minimizer is then obtained solving the associated Euler-Lagrange equations using the Conjugated Gradient (CG) method when the SVD. When available, we use the rSVD of the preprocessing to show that the method works, it is however not necessary and we point this out in Experiment 3.
@@
The DIP has been used in MPI [14]. It is, to our knowledge, the only machine learning approach which has been qualitatively and quantitatively evaluated on the Open MPI dataset. It is based on the idea to drop the explicit regularizer $\mathcal{R}$ in Eq. (2) and to use the implicit prior captured by a neural network parametrization, i.e., if $\varphi_{\theta}$ is a neural network with parameters $\theta$ taking as input a noisy array $z$, we can train the network $\varphi_{\theta}$ by solving the following minimization problem
@@
to obtain a reconstruction $u=\varphi_{\theta^{\ast}}(z)$. For details on our implementation we refer to Sec. S3 in the supplementary material.

### v2

@@ 2.3 ZeroShot-PnP: The Proposed Plug-and-Play Approach with a Zero-Shot Denoiser
Choosing auxiliary variables $u_{1}$, $u_{2}$ and $u_{3}$ we decouple the quantities in Equation (3), in particular, we choose the following equivalent constrained minimization problem in consensus form:
@@
At each iteration $k$, the term in Equation (6) is in charge of the regularized inversion of the system matrix whereas the term in Equation (7) performs regularized Gaussian denoising with positivity constraint. Finally, the term in Equation (8) is the proximal mapping of the $\ell^{1}$-norm, i.e., it results in a soft-thresholding of parameter $\alpha/\mu_{k}$.
@@
The idea behind the PnP approach is to substitute the Gaussian denoising in Equation (7) by any other denoiser with positive output and w.r.t. noise level $\lambda/\mu_{k}$, yielding the following algorithm:
@@
The choice of a suitable denoiser in Equation (10) is key because the performance of the algorithm significantly depends on it. In Computer Vision, machine learning based 2D denoisers have gained a lot of interest in recent years. They are typically trained on large datasets and possess a higher generalization ability than smaller models.
The denoiser we employ in this work is the benchmark deep denoiser prior [70]. It is based on a deep CNN architecture (DRUNet) which combines a U-Net [49] and ResNet [28]. The denoiser is publicly available and trained on a large dataset that combines various datasets including BSD [11], Waterloo Exploration Database [44], DIV2K [1] and Flick2K [43]. We do not retrain the denoiser on MPI or MPI-related data, but employ it as a zero-shot denoiser (in the sense that it has not seen the domain before).
We perform the denoising of the 3D target in a slicewise fashion strategy [4], i.e., at each iteration $k$ we consider the 2D slices perpendicular to an axis and perform the denoising of each 2D slice with the 2D denoiser. This slicewise denoising is performed at step $k$ along each axis and the three resulting volumes are averaged. The overall algorithm is given in Algorithm 1.
We remark here that if we consider no $\ell^{1}$-prior the same HQS yields the following iterative scheme, which can be regarded as a special case of Algorithm 1, which we will call ZeroShot-PnP in Section 3:
$\begin{split}u_{1}^{k+1}&=\arg\min_{u_{1}}\lVert f-Au_{1}\rVert_{2}^{2}+\mu_{k}\left\lVert u_{1}-u_{2}^{k}\right\rVert_{2}^{2}\\ u_{2}^{k+1}&=\mathrm{Denoiser}\left(u_{1}^{k+1}\,,\sqrt{\lambda/\mu_{k}}\right).\end{split}$ (12)
In Section 3 we will see that both the ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP have benefits, especially on the Open MPI dataset.
Algorithm 1 The proposed ZeroShot-$\ell^{1}$-PnP algorithm.
Input: scan data $f$ and system matrix $A$, parameters $\alpha$, $n_{\mathrm{it}}$, $\mu_{0}$.
@@
1:$u_{2}^{0},u_{3}^{0}\leftarrow 0$;
2:$k\leftarrow 0$;
3:while $k\leq n_{\mathrm{it}}$ do
4:     $u_{1}^{k+1}\leftarrow$ConjGrad$\left(A^{\ast}A+\mu_{k}\mathrm{Id}\,,A^{\ast}f+\mu_{k}\frac{u_{2}^{k}+u_{3}^{k}}{2}\right)$;
5:     $\sigma_{k}\leftarrow\sqrt{\mathrm{Var}(u_{1}^{k+1})}$;
6:     if $k=0$ then
7:         $\lambda\leftarrow\mu_{0}\cdot\sigma_{0}^{2}$
8:     end if
9:     $u_{2}^{k+1}\leftarrow$Slicewise-Denoise$\left(u_{1}^{k+1}\,,\sigma_{k}\right)$; $\triangleright$ ZeroShot-Denoiser
10:     $u_{3}^{k+1}\leftarrow$Soft-Threshold$\left(u_{1}^{k+1}\,,\frac{\alpha}{\mu_{k}}\right)$;
11:     $\mu_{k+1}\leftarrow\lambda/\sigma_{k}^{2}$;
12:     $k\leftarrow k+1$;
13:end whilereturn $u_{2}^{k+1}$
@@ 2.4 Parameter Selection Strategies
The ZeroShot-$\ell^{1}$-PnP approach proposed depends on the (hyper)parameters $\mu_{k}$, $\lambda$ and $\alpha$ as well as on the number of iterations $k$ of the scheme (stopping criterion). We remark that the $\sigma_{k}$ are not free parameters, but related via $\sigma_{k}=\sqrt{\lambda/\mu_{k}}$. Typically, in PnP approaches (e.g., [70]) the corresponding parameters are set fixed, in particular the noise related parameter $\sigma_{k}$ fed into the denoiser is set fixed for each iteration $k$ for a fixed number of iterations $n_{\mathrm{it}}$. We now propose more adaptive ways to choose the above parameters.
@@
where $\bar{u}_{1}^{k}$ is the mean of the vector $u_{1}^{k}$. The value $\tilde{\sigma}_{k}^{2}$ is an upper bound on the noise level (and potentially can be seen as a very coarse estimate which empirically turns out to be sufficient for our purposes; another interpretation can be as a normalization). Then, we set the $\mu_{k}$ by
@@
To determine $\lambda=\tilde{\lambda},$ we use the relation $\tilde{\lambda}=\tilde{\mu}_{0}\tilde{\sigma}_{0}^{2},$ where we define $\tilde{\sigma}_{0}^{2}$ via Equation (13), and $\tilde{\mu}_{0}$ by the following observation: initializing $u_{2}^{0}=u_{3}^{0}=0$ in Algorithm 1, $u_{1}^{0}$ is the solution of the Tikhonov regularization problem (cf. Equation (9) and Equation (2) with $\mathcal{R}(u)=\lVert u\rVert_{2}^{2}$). Consequently, one can set the parameter $\mu_{0}$ to
@@
where $\tilde{\lambda}_{\mathrm{Tik}}$ denotes an estimate on the parameter $\lambda_{\mathrm{Tik}}$ of the corresponding Tikhonov problem.
Finally, the algorithm is an iterative scheme that stops after $n_{\mathrm{it}}$ iterations and whose parameters are automatically updated at each iteration by the estimated parameter $\tilde{\sigma}_{k}$ in Equation (13). The only parameters that have to be provided are $\alpha$, $n_{\mathrm{it}}$ and the starting Tikhonov parameter $\mu_{0}=\tilde{\lambda}_{\mathrm{Tik}}$. We have validated $n_{\mathrm{it}}$ and $\mu_{0}$ on a simulated data set which we describe in Section 2.5.
@@
In each iteration a soft-thresholding (proximal mapping of the $\ell^{1}$ norm) of parameter $\alpha/\mu_{k}$ is performed (cf. Equation (11)), and thus all values of magnitude below $\alpha/\mu_{k}$ are set to zero. Let $u$ be the ground truth particle distribution and $u_{\mathrm{rec}}$ its reconstruction using the system matrix $A$. The matrix $A$ is obtained by scanning a delta distribution with a known concentration of $c_{\delta}$ mmol. It follows that if $c_{u}$ is the maximum concentration level of the ground truth $u$, the $\ell^{\infty}$-norm of the reconstruction is bounded from above $\lVert u_{\mathrm{rec}}\rVert_{\infty}\leq c_{u}/c_{\delta}=:c_{\mathrm{rec}}$. Theqrefore, to avoid the all zeros reconstruction, it is necessary that $\alpha/\mu_{k}<c_{\mathrm{rec}}$, i.e., that $\alpha<c_{\mathrm{rec}}\cdot\mu_{0}$. In particular, if some knowledge about the concentration injected and the concentration of the delta distribution is available, it is possible to set a rough bound on the parameter $\alpha$. In the experiments in Section 3, because in the Open MPI Dataset $c_{\delta}=100$mmol and we know that for each phantom $c_{u}\leq 100$mmol, we can deduce that $c_{\mathrm{rec}}\leq 1$ and that $\alpha\leq\mu_{0}$. Theqrefore, it is reasonable to choose $\alpha$ as some percentage of $\mu_{0}$. In this work we have set $\alpha=0.005\mu_{0}$ for all experiments.
2.5 Hybrid Dataset for Parameter Validation
For the validation of the parameters of both the proposed methods and the baseline methods we have produced a hybrid MPI dataset. In more detail, we have produced concentrations $u_{\mathrm{GT}}$ on a $19\times 19\times 19$ grid that play the role of MPI phantoms. We have considered the system matrix $A$ provided in the Open MPI dataset and without preprocessing we have applied it to the ground truths $u_{\mathrm{GT}}$ to obtain a signal $f$ with realistic noise distribution (since $A$ is composed of real noisy delta scans) and added an additional Gaussian noise $\eta$ obtaining a dataset of scans of the form $f=Au_{\mathrm{GT}}+\eta$. The total dataset consists of 30 ground truths, generated to provide the same challenges for which the Open MPI dataset has been created, i.e., to test the reconstruction capabilities w.r.t. shape, different concentration levels and resolution. In particular, we have produced 10 cone-shaped phantoms, 10 graph-like phantoms and 10 additional phantoms composed of small dot-like regions with different concentration levels. The cones where produced with random sizes, shifts and orientations inside the grid. The graph like-phantoms were generated with the following procedure: a number of vertices $V$ between 4 and 6 is randomly chosen and $V-1$ distinct edges (pairs of vertices) are randomly selected; the voxels thorough which an edge passes or that contain a vertex get value $1$ whereas every other voxel stays zero. The resulting phantom is then blurred with a 3D normal Gaussian filter of variance 1 and thresholded to thicken the edges. Finally, the last set of phantoms is generated with the same procedure as for the graph-like phantoms but considering a randomly selected number of vertices between $6$ and $9$ but without edges. Each vertex gets a concentration level randomly selected between $0.05$ and $1$ and the Gaussian blurring and subsequent tresholding is applied to obtain regions bigger than a single voxel. All phantoms upon generation have maximum value 1 and have been rescaled to a maximum value $\beta\sim U(0.5\,,1.5)$.
We point out that this validation set is all we need to validate the parameters before testing on the Open MPI dataset and that no training set is necessary because no training is performed: we use a general purpose denoiser [70] pre-trained on natural images and we do not train and ad hoc denoiser.
@@ 3 Results
In this section we quantitatively and qualitatively evaluate both the ZeroShot-$\ell^{1}$-PnP in Algorithm 1 and the ZeroShot-PnP variant in Equation (12) in Section 2.3 on the OpenMPI dataset [39]. Qualitative comparision using the well established quality measures PSNR and SSIM is possible by inferring the ground truth from a CAD model; details are given in Section 3.1. We compare the ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP schemes with the Deep Image Prior (DIP) and with the standard Tikhonov $\ell^{2}$-regularization as a baseline; moreover, we perform additional comparison with the PP-MPI algorithm [4] which uses a denoiser trained on MPI-friendly data. We show the performance of the PP-MPI denoiser when it is used in place of the ZeroShot-Denoiser in Algorihtm 1 (line 9).
Finally, we also show the performance of the PP-MPI denoiser in the ZeroShot-PnP algorithms. All these methods are described in Section 3.2. The evaluation is conducted in Section 3.3.
In particular, in Experiment 1 we compare the results obtained with the DIP, Tikhonov and PP-MPI methods on the 3D Open MPI Dataset using typical preprocessing and validated parameters. In Experiment 2 we focus more on the differences between the proposed ZeroShot-PnP method and the previous PP-MPI algorithm and the different performances of the denoisers employed. In Experiment 3 we test the ZeroShot-PnP with and without the $\ell^{1}$-prior with decreasing levels of preprocessing. Finally, in Experiment 4 we show that MPI reconstructions with the proposed methods are possible with minimal levels of preprocessing and without the employment of the SVD, which becomes more and more cumbersome to compute as the dimensionality of the problem increases.
The preprocessing was implemented both in Matlab and in Python while the ZeroShot-$\ell^{1}$-PnP and ZeroShot-PnP algorithms as well as the baseline algorithms for comparison were implemented in Python 3.9, using Numpy and PyTorch. The experiments were performed on a workstation with 13th Gen Intel(R) Core(TM) i9-13900KS, 128 GB of RAM, an NVIDIA RTX A6000 GPU and Windows 11 Pro.
@@ 3.1 Image Quality Measures
In order to quantitatively assess the reconstruction quality, the Open MPI dataset provides CAD models of the phantoms. This allows to compare the reconstructions with the corresponding ground truth implied by the CAD model. We use both Peak Signal-to-Noise Ratio (PSNR) and the Structural Similarity Index Measure (SSIM) [64] as measures.
@@
For the $\mathrm{SSIM}$ measure we employ the following definition and choice of parameters: given two images $f,g$, the SSIM is defined as $\mathrm{SSIM}(f,g)=\left[l(f,g)\right]^{\alpha}\cdot\left[c(f,g)\right]^{\beta}\cdot\left[s(f,g)\right]^{\gamma}$, where $l(f,g)=\frac{2\mu_{f}\mu_{g}+C_{1}}{\mu_{x}^{2}+\mu_{y}^{2}+C_{1}}$, $c(f,g)=\frac{2\sigma_{f}\sigma_{g}+C_{2}}{\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2}}$, $s(f,g)=\frac{\sigma_{fg}+C_{3}}{\sigma_{f}\sigma_{g}+C_{3}}$ are the luminance (l), contrast (c) and structure (s)  [64] and $\mu_{f}$ is the mean value, $\sigma_{f}$ the standard deviation of $f$; $\mu_{g}$ and $\sigma_{g}$ are analogously defined for $g$ and $\sigma_{fg}$ the covariance of $f$ and $g$. Following the implementation of [14], we set $\alpha=\beta=\gamma=1$ and $C_{1}=(0.01\cdot R)^{2}$, $C_{2}=(0.03\cdot R)^{2}$ and $C_{3}=0.5\cdot C_{2}$ where $R=100$, i.e., the concentration of the Delta Sample used during the calibration process. In what follows we always display $\mathrm{PSNR}_{\mathrm{max}}$ and $\mathrm{SSIM}_{\mathrm{max}}$ when showing the reconstruction of the Open MPI dataset.
@@ 3.2 Methods for Comparison
The inversion of the system matrix is performed considering the minimization problem in Equation  (2) with $\mathcal{R}(u)=\lVert u\rVert_{2}^{2}$; the minimizer is then obtained solving the associated Euler-Lagrange equations using the Conjugated Gradient (CG) method when the SVD is not available. When the SVD is indeed available, we use it for faster direct solutions of the inversion problems, it is however not necessary as we point out in Experiment 4.
@@
The DIP has been used in MPI [14]. It is, to our knowledge, the only machine learning approach which has been both qualitatively and quantitatively evaluated on the Open MPI dataset. It is based on the idea of dropping the explicit regularizer $\mathcal{R}$ in Equation  (2) and of using the implicit prior captured by a neural network parametrization, i.e., if $\varphi_{\theta}$ is a neural network with parameters $\theta$ taking as input a noisy array $z$, we can train the network $\varphi_{\theta}$ by solving the following minimization problem
@@
[…]
The PP-MPI Algorithm with Trained Denoiser
We now briefly recall the PP-MPI algorithm in [4]. The authors consider the following equivalent formulation of the regularized inversion of the system matrix $A$:
$\arg\min_{u}\sum_{i}\alpha_{i}g_{i}(u)\quad\text{s.t.}\quad\lVert Au-f\rVert_{2}<\epsilon$ (19)
where $\alpha_{i}$ is the parameter weighting the i-th regularization function $g_{i}$, the data fidelity term for given data $f$ and is imposed as a constraint; $\epsilon$ is the upper bound on the noise level. The constrained minimization problem in Equation (19) can be solved using an ADMM splitting scheme that yields the following algorithm:
$\begin{split}u^{n+1}&=\left(\mathbb{I}+A^{*}A\right)^{-1}\left(A^{*}(z_{0}^{n}+d_{0}^{n})+z_{1}^{n}+d_{1}^{n}\right)\\ z_{1}^{n+1}&=\mathrm{Denoiser}(u^{n+1}-d_{1}^{n})\\ d_{1}^{n+1}&=d_{1}^{n}+z_{1}^{n+1}-u^{n+1}\\ z_{0}^{n+1}&=\Psi_{f,\epsilon}(Au^{n+1}-d_{0}^{n})\\ d_{0}^{n+1}&=d_{0}^{n}+z_{0}^{n+1}-Au^{n+1}\\ \end{split}$ (20)
where $\Psi_{f,\epsilon}$ is the projection function onto the ball of radius $\epsilon$ centered at $f$, taking care of the data fidelity constraint. The denoiser trained by the authors [4] on MPI-friendly data will be called PP-MPI-Denoiser in what follows.